A random variable X has the following probability function:
Values of X, x: 0 1 2 3 4 5 6 7
p(x): 0 k 2k 2k 3k k2 2k2 7k2+ k


(i) Find k, (ii) Evaluate P(X < 6), (P X ≥ ),6 and 0(P < X < ),5 (iii) If ,
2
1
(P X ≤ )a >
Find the minimum value of a, and (iv) Determine the distribution function of X.
1
Expert's answer
2020-03-16T13:07:02-0400

A random variable "X" has the following probability function:  


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Values\ of\ x & 0 & 1 & 2 & 3 & 4 & 5 & 6 & 7 \\ \hline
 P(X=x) & 0 & k & 2k & 2k & 3k & k^2 & 2k^2 & 7k^2+k
\end{array}"

(i) Find "k"

For a discrete probability distribution of a random variable  "\displaystyle\sum_{i=1}^nP(X=x_i)=1." Then


"0+k+2k+2k+3k+k^2+2k^2+7k^2+k=1,\ k\geq0""10k^2+9k-1=0""(k+1)(10k-1)=0"

Since "k\geq0," we take "k=\dfrac{1}{10}."


(ii) Evaluate "P(X<6),\ P(X\geq6)" and "P(0<X<5)"

"P(X<6)=P(X=0)+P(X=1)+P(X=2)+""+P(X=3)+P(X=4)+P(X=5)=""1-P(X=6)-P(X=7)=""=1-2(0.1)^2-(7(0.1)^2+0.1)=0.81"

"P(X\geq6)=P(X=6)+P(X=7)=""=2(0.1)^2(7(0.1)^2+0.1)=0.19"


"P(0<X<5)=P(X=1)+P(X=2)+""+P(X=3)+P(X=4)=""=0.1+2(0,1)+2(0.1)+3(0.1)=0.8"

(iii) If "P(X\leq a)>{1 \over 2}," find the minimum value of "a"


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Values\ of\ x & 0 & 1 & 2 & 3 & 4 & 5 & 6 & 7 \\ \hline
 P(X=x) & 0 & 0.1 & 0.2 & 0.2 & 0.3 & 0.01 & 0.02 & 0.17
\end{array}"

"P(X\leq3)=0+0.1+0.2+0.2=0.5"

"P(X\leq4)=0+0.1+0.2+0.2+0.3=0.8>0.5"

The minimum value of "a" is "4."


(iv) Determine the distribution function of "X."

"F(0)=P(X\leq0)=0"


"F(1)=P(X\leq1)=0+k=k=0.1"


"F(2)=P(X\leq2)=0+k+2k=3k=0.3"


"F(3)=P(X\leq3)=0+k+2k+2k=5k=0.5"


"F(4)=P(X\leq4)=0+k+2k+2k+3k=8k=0.8"


"F(5)=P(X\leq5)=0+k+2k+2k+3k+k^2="

"=8k+k^2=0.81"


"F(6)=P(X\leq6)=0+k+2k+2k+3k+k^2+2k^2="

"=8k+3k^2=0.83"


"F(7)=P(X\leq7)=10k^2+9k=1"



"F(x)=x = \begin{cases}
 0 &\text{ } x<1 \\
 0.1 &\text{ } x<2 \\
 0.3 &\text{ } x<3 \\
 0.5 &\text{ } x<4 \\
 0.8 &\text{ } x<5 \\
 0.81 &\text{ } x<6 \\
 0.83 &\text{ } x<7 \\
 1 &\text{ } x\geq7
\end{cases}"